# Instance Segmentation of Scene Sketches Using Natural Image Priors

### v1

@@ Abstract
Sketch segmentation involves grouping pixels within a sketch that belong to the same object or instance. It serves as a valuable tool for sketch editing tasks, such as moving, scaling, or removing specific components. While image segmentation models have demonstrated remarkable capabilities in recent years, sketches present unique challenges for these models due to their sparse nature and wide variation in styles. We introduce SketchSeg, a method for instance segmentation of raster scene sketches. Our approach adapts state-of-the-art image segmentation and object detection models to the sketch domain by employing class-agnostic fine-tuning and refining segmentation masks using depth cues. Furthermore, our method organizes sketches into sorted layers, where occluded instances are inpainted, enabling advanced sketch editing applications. As existing datasets in this domain lack variation in sketch styles, we construct a synthetic scene sketch segmentation dataset featuring sketches with diverse brush strokes and varying levels of detail. We use this dataset to demonstrate the robustness of our approach and will release it to promote further research in the field.
Figure 1: Our method performs instance segmentation of raster sketches. It effectively handles diverse types of sketches, accommodating variations in stroke style and complexity.
1 Introduction
@@
Sketches serve as a powerful tool for visual exploration, ideation, and planning. Traditional sketching workflows often begin with artists working on a single canvas layer (either physical or digital) to maintain creative momentum. As the sketch evolves and requires refinement (e.g., adjustments to composition, perspective, or other elements), artists face the tedious task of manually segmenting different elements of the sketch into discrete, editable layers. Automating the sketch segmentation process offers a promising solution. However, this task presents unique challenges due to the sparse and abstract nature of line drawings, as well as the inherent variability in human sketching styles. Existing methods for scene-level sketch segmentation typically rely on training dedicated models using annotated sketch datasets [26, 47, 5]. However, most available datasets are confined to specific sketch styles and a limited set of object categories, restricting the generalization capabilities of existing methods.
@@
In this work, we introduce SketchSeg, a method for instance segmentation of raster scene sketches that outperforms previous approaches in accommodating a wider variety of sketch styles and concepts. We use the segmentation map to divide the sketch into sorted layers to support effective sketch editing.
@@
Our method builds upon Grounded SAM [24], a state-of-the-art approach for open-vocabulary image segmentation, which has demonstrated remarkable capabilities in segmenting complex scenes across diverse object categories. Grounded SAM combines two models to achieve this: Grounding DINO [17] for object detection and Segment Anything (SAM) [12] for mask generation. We analyze the performance of these models on sketch inputs, revealing that the domain gap between real and sketched objects presents significant challenges for Grounding DINO. In contrast, SAM exhibits a surprising ability to generalize to sketches, though it still faces difficulties specific to the sketch domain. To address the gap in object detection, we fine-tune Grounding DINO on a small subset of annotated scene sketches from the SketchyScene dataset [47]. Our straightforward fine-tuning technique proves highly effective, achieving a substantial improvement in Grounding DINO’s detection performance on sketches, with Average Precision increasing from 24% to 75%.
@@
For object segmentation, we apply SAM [12] in the sketch domain, using detected object regions from our finetuned Grounding DINO. This is followed by a depth-based refinement stage to resolve ambiguities in overlapping regions. Finally, we decompose the segmented sketch into sorted layers and employ a pretrained image inpainting model [31] to fill in missing regions. This layered representation facilitates sketch editing, allowing users to drag or manipulate segmented objects without the need to manually sketch the affected regions, as we demonstrate in the provided video.
@@
To evaluate our method on diverse scene sketches, we construct a synthetic annotated dataset that extends existing benchmarks along three key dimensions: drawing style, stroke style, and object categories. The dataset integrates two complementary pipelines to enhance diversity. The first pipeline builds on SketchyScene [47], expanding its clipart-like sketches with styles ranging from high-fidelity representations to symbolic, abstract sketches, introducing challenging out-of-distribution cases. The scenes are created in vector format to allow for stroke style variations, including Calligraphic Pen, Charcoal, and Brush Pen styles. The second pipeline leverages the Visual Genome dataset [13], which provides annotated scenes with object variety. Using the InstantStyle method [33], we generate expressive, natural-looking sketches spanning 74 categories, extending SketchyScene’s original 45 categories by 54 new categories. Our dataset contains 20,000 annotated scene sketches in total, and is highly extensible. Our evaluations demonstrate that SketchSeg generalizes well to these challenging variations, significantly advancing the state of the art.
@@
Figure 2: Overview of the sketch segmentation pipeline. Given an input sketch image, our framework first detects bounding boxes using a customized Grounding DINO to obtain region proposals, and then perform segmentation with SAM models. The localization and segmentation are refined by incorporating the depth features. The result segmentation can be viewed as a layered decomposition of object components in the original sketch.
2 Related Work
2.1 Part-Level Sketch Segmentation
@@
The majority of work in the sketch segmentation domain focuses on part-level semantic segmentation, in which the goal is to assign labels to object parts (e.g., the body, wings, and head of a bird). These methods often rely on curated part-level sketch segmentation datasets [8, 14, 36, 6, 10] to train a segmentation model, and use various network architectures, including CNNs [46, 32], RNNs [22, 36, 11], Graph Neural Networks [44, 40], Transformers [34, 45], and more specific techniques such as deformation networks [20] and CRFs [27]. These approaches typically operate on a fixed set of object classes, and recognize a predefined set of object parts within them. Other work focuses on perceptual grouping [14, 15] to achieve class-agnostic segmentation. However all of these methods are designed to tackle part-level segmentation and are not suitable for scene sketches.
2.2 Scene-Level Sketch Segmentation
@@
Scene-level sketch segmentation remains largely under-explored. Qi et al. [21] extend the perceptual grouping approach to scene-level images, forming semantically meaningful groupings of edges, though with limited accuracy on complicated scenes. Zou et al. [47] construct the SketchyScene dataset, providing annotated scene sketches with meaningful layouts of object interactions, and use it to train an instance segmentation model based on the Mask R-CNN architecture [9]. However, their method is limited to the predefined categories included in the dataset, and the proposed dataset contains sketches with clipart-like appearance which challenges the model’s ability to generalize to other artistic styles. Building on SketchyScene, Ge et al. [7] introduced SKY-Scene and TUB-Scene by replacing its object components with sketches from the Sketchy [26] and TU-Berlin [6] datasets. However, their proposed fusion network is fundamentally limited to the fixed set of classes it was trained on, and the trained network weights are not publicly available. SFSD [43] develops a dataset featuring more complex scene sketches, and utilizes a bidirectional LSTM to produce stroke-level segmentation. Unfortunately, the dataset and model are not publicly available. SketchSeger [38] proposes a hierarchical Transformer-based model for semantic sketch segmentation. However their model is inherently restricted to the predefined set of classes used during training. Bourouis et al. [2] finetune the CLIP image encoder [23] on the FS-COCO [5] dataset, leveraging the model’s vision-language prior to enable open-vocabulary scene segmentation. However their method is designed for semantic segmentation, and it struggles to generalize to more challenging sketch styles and scene layouts.
2.3 Image Segmentation
@@
The task of image segmentation have been widely explored [9, 1, 4, 35]. The advent of vision-language models [23, 16, 37] has led to numerous object detection and segmentation methods with impressive generalization capabilities [41, 24, 12, 19]. Grounding DINO [17] is a state-of-the-art object detection model trained on over 10 million images. It builds on top of DINO [3], a strong vision encoder, with effective grounding module that fuses visual and textual information, enabling open-vocabulary detection of unseen objects. Segment Anything (SAM) [12] is an image segmentation model trained on over 11 million images and 1.1 billion masks, capable of producing high-quality object masks based on various forms of conditioning such as bounding boxes. Grounded SAM [24], which our method builds upon, combines Grounding DINO and SAM for open-vocabulary image segmentation, achieving robust performance across diverse object categories. Yet, despite demonstrating impressive capabilities on natural images, we show that these models struggle with segmenting sketches.
3 Method
@@
Given a raster sketch, our goal is to produce a segmentation map such that pixels belonging to the same object instance are grouped together. Based on the segmentation map, we also divide the sketch into layers, sorted by depth. Our pipeline is illustrated in Figure 2. Given the input sketch, we first perform object detection using a fine-tuned Grounding DINO model, which produces a set of candidate object bounding boxes. These bounding boxes are then used to produce an initial set of object masks with a pre-trained Segment Anything (SAM) [12] model. Next, we perform a refinement stage that leverages scene depth information and classical morphological operations to assign the final segmentation. This stage also employs a pre-trained inpainting model [31] to produce scene layers.
3.1 Sketch-Aware Object Detection
@@
Grounding DINO [17] is an object detection model which outputs bounding boxes for recognized object instances, based on a given text prompt describing the scene. While effective for natural images, the model in its original configuration demonstrates limited generalization to sketches (we show this numerically in Section 5). To address this limitation, we fine-tune Grounding DINO on sketches. The largest available annotated sketch dataset containing complex scenes is SketchyScene [47]. It contains 30K segmented sketches across 45 class labels. We find that a naive fine-tuning with the SketchyScene data leads to severe overfitting to the small set of predefined object classes.
@@
To overcome this overfitting, we propose a class-agnostic fine-tuning strategy. Instead of relying on predefined class labels, we train the model to distinguish between instances based on their visual characteristics, aiming to push the model to rely on Gestalt properties such as closure, continuity, and emergence, to group together strokes forming a single object. Specifically, we utilize a small subset of 5000 sketches from the SketchyScene dataset, and consolidate their class labels into a single label, “object”. We use a Grounding DINO model initialized with a pretrained Swin Transformer [18] backbone and fine-tune the model’s detection head for bounding box prediction. For training, we employ standard object detection losses used in the original Grounding DINO training (Focal Loss, L1 Loss, and GIoU Loss), while eliminating the class recognition loss. At inference, the model is prompted with the input image and the word “object” to detect all potential object instances in the scene. This results in an initial set of $k$ bounding boxes $B=\{B_{i}\}_{i=1}^{k}$ and a confidence score per bounding box.
3.2 Mask Extraction and Bounding Boxes Refinement
@@
Once the bounding boxes are obtained, we use a pretrained Segment Anything (SAM) model [12] to extract masks for the corresponding objects directly from the sketch. This results in an initial set of $k$ masks $M=\{M_{i}\}_{i=1}^{k}$ which we refine using simple binary operations such as morphological closing and flood-fill, to eliminate small artifacts.
@@
Next, we use the refined masks to enhance the set of generated bounding boxes. A common practice is to eliminate redundant bounding boxes often corresponding to the same object (such as $B_{i},B_{j}$ shown in red and blue in Figure 2) using Non-Maximum Suppression (NMS), which filters out bounding boxes with low confidence scores that has significant intersection with others. However, IoU of the bounding boxes may not reliably reflect object overlap in cases where objects do not fully cover the pixels in their bounding boxes. This issue is especially pronounced in sketches, which are sparser than photorealistic images. We use the initial set of masks to compute a more fine-grained IoU. Specifically, for a pair of overlapping bounding boxes $B_{i}$ and $B_{j}$, we extract the regions within the bounding boxes that intersect with the sketch: $M_{i}*S,M_{j}*S$, and compute the IoU of these regions to define an “overlapping” score between two objects $i,j$:
$\mathcal{O}(i,j)=\text{IoU}(M_{i}*S,M_{j}*S).$ (1)
@@
Figure 3: Resolving ambiguities in overlapping regions. The depth map $D$ is sampled along sketch pixels at evenly spaced points, and the sampled points are grouped by their corresponding object (e.g., $P_{i}$ corresponds to the $i$’th object). Each object is assigned a depth score based on the majority of depth values from the sampled points. Ambiguous pixels are then assigned to the mask with the highest depth score, prioritizing foreground objects.
The filtered set of masks may still include overlapping regions, as illustrated in Fig. 3, where it is unclear which instance the pixels should be associated with. To resolve such overlaps and assign each pixel to a single object instance, we give priority to objects in the foreground. We utilize DepthAnything [39] to extract the depth map $D$ of the input sketch. We then sample $D$ along the sketch pixels at equally spaced points $P=\{p_{1},p_{2},\dots,p_{n}\}$, analogous to projecting rays through the sketch pixels to the scene. For each mask $M_{i}$, we identify the subset of points that lie within the mask: $P_{i}=\{p\in P|p\in M_{i}\}$. For example, in Figure 3, $P_{i}$ represents the set of points belonging to the sofa, while $P_{j}$ denotes the set of points belonging to the cat. For each point $p\in P_{i}$ we associate a depth value $D(p)$ using the depth map. We then compute a depth score for each mask as the mode of the depth values associated with its sampled points:
$ds(M_{i})=\arg\max_{D(p)}\text{count}(\{D(p)|p\in P_{i}\}).$ (2)
Based on this score, we assign ambiguous pixels to the mask with the highest depth score, ensuring that foreground objects take precedence. Lastly, to ensure complete coverage of the sketch, we employ a watershed-based [28] refinement, propagating existing mask labels to previously unlabeled sketch pixels.
3.3 Layer Inpainting
As a final step we extract complete layers for each object in the sketch, inpainting any occluded regions using a pretrained SDXL inpainting model [31]. The goal of this stage is to support basic sketch editing operations, such as translation and scaling. We isolate each object $i$ by intersecting the sketch with its corresponding mask $M_{i}*S$ (Figure 13). We then identify the group of masks that intersect with $M_{i}$: $\mathcal{H}(M_{i})=\{M_{j}|M_{j}\cap M_{i}\neq\varnothing\}$. Finally, we define the inpainting mask $C_{i}$ as the intersection of $\mathcal{H}(M_{i})$ with the object’s bounding box: $C_{i}=\mathcal{H}(M_{i})\cap B_{i}$ (shown in green in Fig. 13), and feed it into the pretrained inpainting model.
3.4 Implementation Details

### v2

@@ Abstract.
Sketch segmentation involves grouping pixels within a sketch that belong to the same object or instance. It serves as a valuable tool for sketch editing tasks, such as moving, scaling, or removing specific components. While image segmentation models have demonstrated remarkable capabilities in recent years, sketches present unique challenges for these models due to their sparse nature and wide variation in styles. We introduce InkLayer, a method for instance segmentation of raster scene sketches. Our approach adapts state-of-the-art image segmentation and object detection models to the sketch domain by employing class-agnostic fine-tuning and refining segmentation masks using depth cues. Furthermore, our method organizes sketches into sorted layers, where occluded instances are inpainted, enabling advanced sketch editing applications. As existing datasets in this domain lack variation in sketch styles, we construct a synthetic scene sketch segmentation dataset, InkScenes, featuring sketches with diverse brush strokes and varying levels of detail. We use this dataset to demonstrate the robustness of our approach. Code and data for this paper are released at project page: https://inklayer.github.io.
Sketch Segmentation, Sketch Understanding, Sketch Editing
submissionid: 224journalyear: 2025conference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ; August 10–14, 2025; Vancouver, BC, Canadabooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers (SIGGRAPH Conference Papers ’25), August 10–14, 2025, Vancouver, BC, Canadadoi: 10.1145/3721238.3730606isbn: 979-8-4007-1540-2/2025/08ccs: Computing methodologies Image segmentationccs: Computing methodologies Non-photorealistic renderingccs: Computing methodologies Scene understanding
1. Introduction
@@
Sketches serve as a powerful tool for visual exploration, ideation, and planning. Traditional sketching workflows often begin with artists working on a single canvas layer (either physical or digital) to maintain creative momentum. As the sketch evolves and requires refinement (e.g., adjustments to composition, perspective, or other elements), artists face the tedious task of manually segmenting different elements of the sketch into discrete, editable layers. Automating the sketch segmentation process offers a promising solution. However, this task presents unique challenges due to the sparse and abstract nature of line drawings, as well as the inherent variability in human sketching styles. Existing methods for scene-level sketch segmentation typically rely on training dedicated models using annotated sketch datasets (Sangkloy et al., 2016; Zou et al., 2018; Chowdhury et al., 2022). However, most available datasets are confined to specific sketch styles and a limited set of object categories, restricting the generalization capabilities of existing methods.
@@
In this work, we introduce InkLayer, a method for instance segmentation of raster scene sketches that outperforms previous approaches in accommodating a wider variety of sketch styles and concepts. We use the segmentation map to divide the sketch into sorted layers to support effective sketch editing.
@@
Our method builds upon Grounded SAM (Ren et al., 2024), a state-of-the-art approach for open-vocabulary image segmentation, which has demonstrated remarkable capabilities in segmenting complex scenes across diverse object categories. Grounded SAM combines two models to achieve this: Grounding DINO (Liu et al., 2023b) for object detection and Segment Anything (SAM) (Kirillov et al., 2023) for mask generation. We analyze the performance of these models on sketch, revealing that the domain gap between real and sketched objects presents significant challenges for Grounding DINO. In contrast, SAM exhibits a surprising ability to generalize to sketches, though it still faces difficulties specific to the sketch domain. To address the gap in object detection, we fine-tune Grounding DINO on a small subset of annotated scene sketches from the SketchyScene dataset (2018). Our fine-tuning technique achieves a substantial improvement in Grounding DINO’s detection performance on sketches, with Average Precision increasing from 26% to 74%.
@@
For object segmentation, we apply SAM (Kirillov et al., 2023) in the sketch domain, using detected object regions from our finetuned Grounding DINO. This is followed by a depth-based refinement stage to resolve ambiguities in overlapping regions. Finally, we decompose the segmented sketch into sorted layers and employ a pretrained image inpainting model (2022) to fill in missing regions. This layered representation facilitates sketch editing, allowing users to drag or manipulate segmented objects without the need to manually sketch the affected regions, as we demonstrate in the provided video.
@@
To evaluate our method on diverse scene sketches, we construct InkScenes, a synthetic annotated dataset of sketched scenes that extends existing benchmarks along three key dimensions: drawing style, stroke style, and object categories. The dataset integrates two complementary pipelines to enhance diversity. The first pipeline builds on SketchyScene (Zou et al., 2018), expanding its clipart-like sketches with styles ranging from high-fidelity representations to symbolic, abstract sketches, introducing challenging out-of-distribution cases. The scenes are created in vector format to allow for stroke style variations, including Calligraphic Pen, Charcoal, and Brush Pen styles. The second pipeline leverages the Visual Genome dataset (Krishna et al., 2017), which provides annotated scenes with object variety. Using the InstantStyle method (Wang et al., 2024), we generate expressive, natural-looking sketches spanning 72 categories, extending SketchyScene’s original 45 categories by 53 new categories. Our dataset contains 20,542 annotated scene sketches in total, and is highly extensible. Our evaluations demonstrate that InkLayer generalizes well to these challenging variations, significantly advancing the state of the art.
@@
Figure 2. Overview of the sketch segmentation pipeline. Given an input sketch image, our framework first detects bounding boxes using a customized Grounding DINO to obtain region proposals, and then perform segmentation with SAM models. The localization and segmentation are refined by incorporating the depth features. The result segmentation can be viewed as a layered decomposition of object components in the original sketch.
2. Related Work
2.1. Part-Level Sketch Segmentation
@@
The majority of work in the sketch segmentation domain focuses on part-level semantic segmentation, in which the goal is to assign labels to object parts (e.g., the body, wings, and head of a bird). These methods often rely on curated part-level sketch segmentation datasets (Ge et al., 2020; Li et al., 2018; Wu et al., 2018; Eitz et al., 2012; Huang et al., 2014) to train a segmentation model, and use various network architectures, including CNNs (Zhu et al., 2018; Wang et al., 2020), RNNs (Qi and Tan, 2019; Wu et al., 2018; Kaiyrbekov and Sezgin, 2020), Graph Neural Networks (Zheng et al., 2022; Yang et al., 2021), Transformers (Wang and Li, 2024; Zheng et al., 2023), and more specific techniques such as deformation networks (Qi et al., 2021) and CRFs (Schneider and Tuytelaars, 2016). These approaches typically operate on a fixed set of object classes, and recognize a predefined set of object parts within them. Other work focuses on perceptual grouping (Li et al., 2018, 2019, 2022) to achieve class-agnostic segmentation. However these methods are designed to tackle part-level segmentation and are not suitable for scene sketches.
2.2. Scene-Level Sketch Segmentation
@@
Scene-level sketch segmentation remains largely under-explored. Qi et al. (2015) extend the perceptual grouping approach to scene-level images, forming semantically meaningful groupings of edges, though with limited accuracy on complicated scenes. Zou et al. (2018) construct the SketchyScene dataset, providing annotated scene sketches with meaningful layouts of object interactions, and use it to train an instance segmentation model based on the Mask R-CNN architecture (He et al., 2018). However, their method is limited to the predefined categories included in the dataset, and the proposed dataset contains sketches with clipart-like appearance which challenges the model’s ability to generalize to other artistic styles. Building on SketchyScene, Ge et al. (2022) introduced SKY-Scene and TUB-Scene by replacing its object components with sketches from the Sketchy (2016) and TU-Berlin (2012) datasets. However, their proposed fusion network is fundamentally limited to the fixed set of classes it was trained on, and the trained network weights are not publicly available. SFSD (Zhang et al., 2022a) develops a dataset featuring more complex scene sketches, and utilizes a bidirectional LSTM to produce stroke-level segmentation. Unfortunately, the dataset and model are not publicly available. SketchSeger (Yang et al., 2023) proposes a hierarchical Transformer-based model for semantic sketch segmentation. However their model is inherently restricted to the predefined set of classes used during training. Bourouis et al. (2024) finetune the CLIP image encoder (Radford et al., 2021) on the FS-COCO (Chowdhury et al., 2022) dataset, leveraging the model’s vision-language prior to enable open-vocabulary scene segmentation. However their method is designed for semantic segmentation, and it struggles to generalize to more challenging sketch styles and scene layouts.
2.3. Image Segmentation
@@
The task of image segmentation have been widely explored (He et al., 2018; Bolya et al., 2019; Cheng et al., 2022; Wang et al., 2021). The advent of vision-language models (Radford et al., 2021; Liu et al., 2023a; Xiao et al., 2023) has led to numerous object detection and segmentation methods with impressive generalization capabilities (Zhang et al., 2022b; Ren et al., 2024; Kirillov et al., 2023; Minderer et al., 2022). Grounding DINO (Liu et al., 2023b) is a state-of-the-art object detection model trained on over 10 million images. It builds on top of DINO (Caron et al., 2021), a strong vision encoder, with effective grounding module that fuses visual and textual information, enabling open-vocabulary detection of unseen objects. Segment Anything (SAM) (Kirillov et al., 2023) is an image segmentation model trained on over 11 million images and 1.1 billion masks, capable of producing high-quality object masks based on various forms of conditioning such as bounding boxes. Grounded SAM (Ren et al., 2024), which our method builds upon, combines Grounding DINO and SAM for open-vocabulary image segmentation, achieving robust performance across diverse object categories. Yet, despite demonstrating impressive capabilities on natural images, we show that these models struggle with segmenting sketches.
3. Method
@@
Given a raster sketch, our goal is to produce a segmentation map such that pixels belonging to the same object instance are grouped together. Based on the segmentation map, we also divide the sketch into layers, sorted by depth. Our pipeline is illustrated in Figure 2. Given the input sketch, we first perform object detection using a fine-tuned Grounding DINO model, which produces a set of candidate object bounding boxes. These bounding boxes are then used to produce an initial set of object masks with a pre-trained Segment Anything (SAM) (Kirillov et al., 2023) model. Next, we perform a refinement stage that leverages scene depth information to assign the final segmentation. This stage also employs a pre-trained inpainting model (von Platen et al., 2022) to produce scene layers.
3.1. Sketch-Aware Object Detection
@@
Grounding DINO (Liu et al., 2023b) is an object detection model which outputs bounding boxes for recognized object instances, based on a given text prompt describing the scene. While effective for natural images, the model in its original configuration demonstrates limited generalization to sketches (we show this numerically in Section 5). To address this limitation, we fine-tune Grounding DINO on sketches. The largest available annotated sketch dataset containing complex scenes is SketchyScene (Zou et al., 2018). It contains 30K segmented sketches across 45 class labels. We find that a naive fine-tuning with the SketchyScene data leads to severe overfitting to the small set of predefined object classes.
@@
To overcome this overfitting, we propose a class-agnostic fine-tuning strategy. Instead of relying on predefined class labels, we train the model to distinguish between instances based on their visual characteristics, aiming to push the model to rely on Gestalt properties such as closure, continuity, and emergence, to group together strokes forming a single object. Specifically, we utilize a small subset of 5000 sketches from the SketchyScene dataset, and consolidate their class labels into a single label, “object”. We use a Grounding DINO model initialized with a pretrained Swin Transformer (Liu et al., 2021) backbone and fine-tune the model’s detection head for bounding box prediction. For training, we employ standard object detection losses used in the original Grounding DINO training (Focal Loss, L1 Loss, and GIoU Loss), while eliminating the class recognition loss. At inference, the model is prompted with the input image and the word “object” to detect all potential object instances in the scene. This results in an initial set of $k$ bounding boxes $B=\{B_{i}\}_{i=1}^{k}$ and a confidence score per bounding box.
3.2. Mask Extraction and Bounding Boxes Refinement
@@
Once the bounding boxes are obtained, we use a pretrained Segment Anything (SAM) model (Kirillov et al., 2023) to extract masks for the corresponding objects directly from the sketch. This results in an initial set of $k$ masks $M=\{M_{i}\}_{i=1}^{k}$ which we refine using simple binary operations such as morphological closing and flood-fill, to eliminate small artifacts.
@@
Next, we use the refined masks to enhance the set of generated bounding boxes. A common practice is to eliminate redundant bounding boxes often corresponding to the same object (such as $B_{i},B_{j}$ shown in red and blue in Figure 2) using Non-Maximum Suppression (NMS), which filters out bounding boxes with low confidence scores that has significant intersection with others. However, IoU of the bounding boxes may not reliably reflect object overlap in cases where objects do not fully cover the pixels in their bounding boxes. This issue is especially pronounced in sketches, which are sparser than photorealistic images. We use the initial set of masks to compute a more fine-grained IoU. Specifically, for a pair of overlapping bounding boxes $B_{i}$ and $B_{j}$, we extract the regions within the bounding boxes that intersect with the sketch $S$: $M_{i}*S,M_{j}*S$, and compute the IoU of these regions to define an “overlapping” score between two objects $i,j$:
(1) $\mathcal{O}(i,j)=\text{IoU}(M_{i}*S,M_{j}*S).$
@@
Figure 3. Resolving ambiguities in overlapping regions. The depth map $D$ is sampled along sketch pixels at evenly spaced points, and the sampled points are grouped by their corresponding object (e.g., $P_{i}$ corresponds to the $i$’th object). Each object is assigned a depth score based on the majority of depth values from the sampled points. Ambiguous pixels are then assigned to the mask with the highest depth score, prioritizing foreground objects.
The filtered set of masks may still include overlapping regions, as illustrated in Fig. 3, where it is unclear which instance the pixels should be associated with. To resolve such overlaps and assign each pixel to a single object instance, we give priority to objects in the foreground. We utilize DepthAnythingV2 (Yang et al., 2024) to extract the depth map $D$ of the input sketch. We then sample $D$ along the sketch pixels at equally spaced points $P=\{p_{1},p_{2},\dots,p_{n}\}$, analogous to projecting rays through the sketch pixels to the scene. For each mask $M_{i}$, we identify the subset of points that lie within the mask: $P_{i}=\{p\in P|p\in M_{i}\}$. For example, in Figure 3, $P_{i}$ represents the set of points belonging to the sofa, while $P_{j}$ denotes the set of points belonging to the cat. For each point $p\in P_{i}$ we associate a depth value $D(p)$ using the depth map. We then compute a depth score for each mask as the mode of the depth values associated with its sampled points:
(2) $ds(M_{i})=\arg\max_{D(p)}\text{count}(\{D(p)|p\in P_{i}\}).$
Based on this score, we assign ambiguous pixels to the mask with the highest depth score, ensuring that foreground objects take precedence. Lastly, to ensure complete coverage of the sketch, we employ a watershed-based (Vincent and Soille, 1991) refinement, propagating existing mask labels to unlabeled sketch pixels.
3.3. Layer Inpainting
As a final step we extract complete layers for each object in the sketch, inpainting any occluded regions using a pretrained SDXL inpainting model (von Platen et al., 2022). The goal of this stage is to support basic sketch editing operations, such as translation and scaling. We isolate each object $i$ by intersecting the sketch with its corresponding mask $M_{i}*S$ (Figure 4). We then identify the group of masks that intersect with $M_{i}$: $\mathcal{H}(M_{i})=\{M_{j}|M_{j}\cap M_{i}\neq\varnothing\}$. Finally, we define the inpainting mask $C_{i}$ as the intersection of $\mathcal{H}(M_{i})$ with the object’s bounding box: $C_{i}=\mathcal{H}(M_{i})\cap B_{i}$ (shown in green in Fig. 4), and feed it into the pretrained inpainting model.
3.4. Implementation Details